# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v1

@@ 3.2 Semi-Structured Tables
In this section, we will focus on semi-structured table processing, which refers to techniques for understanding, representing, and extracting information from semi-structured tables.
Compared to structured relational data, semi-structured tables exhibit more complex structural organization, characterized by merged cells, hierarchical headers, and nested tables. In order to give a relatively clear definition of semi-structured tables, we summarize the common characteristics and representation formats of these tables as follows for your reference.
(1) Features of Semi-Structured Tables: We observe that many semi-structured tables share a set of recurring properties, which can be summarized into five core characteristics, illustrated in Figure 5. (1) Wrong Index: the indices of rows or columns may be missing, ambiguous, or inconsistent suhara2022annotating . (2) Hierarchical Content: not all rows or columns adhere to the same schema yang2022tableformer (e.g., a “summary” cell may appear under the “ID” column without any logical connection between the index and the content). (3) Merged Cell: a single cell may span multiple rows or columns guo2022trust . (4) Flexible Header Orientation: the relative position of keys and values is not fixed (e.g., keys may appear to the left of or above their associated values, as observed in the “summary” and “Name” fields). (5) Inconsistent Content: cell values are highly flexible and may include words, numbers, or even entire sentences.
@@
(2) Table Representation Forms: The representation of semi-structured table is varied. We have classified and analyzed common formats and analyzed their characteristics as much as possible for your reference. The first thing we consider is whether the table format can support the representation of semi-structured tables. If so, can it fully represent the special structure of all tables? So we put forward “Semi-Structured Table Support” and “Completeness” to explain these two issues. Besides, we also investigated the structural complexity of different tables and proposed “Structural Complexity”, “Schema Complexity” and “Overall Relative Token Length Level”. “Structural Complexity” considers the complexity of the representation of the external structure of the cell, while “Schema Complexity” considers the complexity of the schema corresponding to each cell in the table structure. And the “Overall Relative Token Length Level” describes the length of the token formed by each form when prompting. This takes into account the complexity of the structure and schema, and is divided into levels 1 to 5 for your reference.
Through our research, we can easily find that tables with simple structure or annotations are often difficult to represent complex semi-structured table structures.(e.g.,CSV, Markdown).At the same time, although some forms (e.g.,JSON) have simple structures, they can be represented by complex schemas.
This inherent complexity poses substantial challenges in aligning queries with both the content and the structure of the table in query answering tasks. The absence of efficient tools (e.g., the openpyxl library for spreadsheet manipulation) and robust representation formats (e.g., Excel or HTML) further exacerbates the difficulty of processing such data.
However, with the aid of LLMs, it becomes possible to directly reason over semi-structured tables, interpret their intricate layouts, and bridge the gap between natural language queries and heterogeneous table structures, thereby enabling more accurate and context-aware query answering.
Traditional table processing approaches like TAPAS herzig2020tapas mainly rely on lightweight models (e.g., BERTdevlin2019bert , RoBERTaliu2019roberta ) to analyze the table. These methods rely on specialized embedding mechanisms to capture the two-dimensional structure of tables, including rows, columns, and numerical order. To support tasks such as cell selection or aggregation, they usually require large-scale weakly supervised pre-training yang2022tableformer . However, due to the small number of parameters and weak NLP capabilities of lightweight models, this method cannot handle some tasks based on semi-structured tables with confusing layouts.
Leveraging the strong semantic understanding capabilities of LLMs, LLM-based approaches generally preprocess the semi-structured table and then perform table understanding or reasoning. We conducted some research on this, and the work is shown in Figure 6.
Figure 6: LLM for Markup Language Overview
$\bullet$
@@
(1) Tree-based Table Modeling: , TUTA wang2021tuta introduces a tree-based encoding framework for semi-structured tables. The root node branches into two child nodes, representing the row and column indices, which are further expanded into hierarchical trees. Each table entry is modeled as a node whose parents are the corresponding row and column nodes. For example, in a personnel information form, the column tree may contain nodes such as “ID” or “Phone Number”, while the row tree represents individual names. ST-Raptor tang2025st proposes a HO-Tree structure to comprehensively encode semi-structured tables. The HO-Tree consists of a Meta-Tree, which captures both the structural and semantic information of table headers, and a Body-Tree, which represents the structure and content of the table body. To construct the HO-Tree, the framework leverages a VLM to detect meta-information, applies heuristic rules to segment tables into subtables, and then builds the HO-Tree using a depth-first search strategy.
@@
(2) Model-based Structuring: TabFormer 10.1145/3726302.3730071 introduces a framework for converting semi-structured tables into relational data. Its chain-of-thought (CoT) module first prompts LLMs to perform step-by-step transformations using “soft operators,” which are table manipulation tools provided through prompts (e.g., split operator to devide the table into subtables if multiple groups of headers are detected). After preprocessing, TabFormer leverages SQL for table reconstruction. Specifically, it first predicts the table indices, then generates the CREATE statements to define the table schema, and finally inserts the extracted contents from the pre-processed tables into the reconstructed relational tables. Auto-Tables li2023autotablessynthesizingmultisteptransformations integrates OCR for structure recognition and applies domain-specific language (DSL) operators. These operators abstract common spreadsheet operations (e.g., melt, transpose), thereby enhancing generalization and adaptability in table conversion tasks. TradeSweep lee2024can , while closely related to Auto-Tables, differs in its approach. Instead of predefined DSL operators, it accepts user-specified transformation requirements and directly generates table conversion code by leveraging a code repository, offering greater flexibility in addressing diverse user needs.
$\bullet$
Table Prompting. Table Prompting means converting table data into LLM’s text input format. The traditional way of table prompting is more like table lineriazation, which simply flattens the contents of all cells into a single string. This crude method will lead to a series of problems such as the prompt sequence being too long and the loss of structured information. Therefore, when using LLM to parse tables, we need to compress the structure and content information of the table. This can not only process large-scale text, but also effectively ensure that the loss of structural information is within a controllable range.
@@
SHEETCOMPRESSOR dong2024spreadsheetllm introduces an efficient framework for serializing large-scale spreadsheets for LLM prompting. Specifically, Structural Anchors identify heterogeneous rows and columns that provide key layout cues, pruning redundant homogeneous ones to yield a skeletonized spreadsheet. Inverted-Index Translation replaces row/column serialization with a lossless JSON dictionary that indexes non-empty cell texts and merges identical values, thereby reducing storage overhead. Finally, Data Format Aggregation clusters adjacent numerical cells of the same type or format, capturing distribution patterns without redundancy. HySem pp2024hysem presents a lightweight LLM pipeline tailored for local deployment. Its Context Optimizer rewrites HTML table tokens into an extremely compressed form, ensuring each cell is represented by only one or two tokens (e.g., “Theme 3: Factors affecting stuff response” is compressed to “Theme 3”). This method establishes a direct mapping between markup and content by minimizing cell representations.
$\bullet$
Table Querying. LLMs typically require an effective and comprehensive method to extract table contents, a process often referred to as table querying. While this process is conceptually similar to SQL, in this section we focus only on approaches that extract information directly from semi-structured data, rather than traditional SQL-based querying.
CoSdong2024spreadsheetllm is a downstream table query program based on SHEETCOMPRESSORdong2024spreadsheetllm , which is mentioned above.First, given a compressed spreadsheet representation and a natural language query, the model performs Table Identification and Boundary Detection, locating the precise subtable relevant to the task. Second, the extracted subtable and query are passed back to the LLM for response generation, enabling focused and context-aware reasoning.
In ST-Raptortang2025st , the table query is based on its HO-Tree. A complex question is first decomposed into sub-questions, each mapped to a sequence of Basic Operations of Ho Tree (e.g.,Retrieval, Condition). To ensure these operations can be reliably executed, the system introduces an Alignment Operation, grounding abstract query terms to actual table headers or values. During execution, the model further leverages Top-down and Bottom-up retrieval strategies to flexibly navigate the hierarchical structure of HO-Tree. For large-scale tables, Column-Type Annotation (discrete/continuous/non-structural) is employed to constrain the search space. Finally, a dual validation mechanism is applied: Forward Constraints check the consistency of intermediate results, while Backward Reasoning reconstructs the question from the generated answer to reduce hallucinations.
3.3 Semi-Structured Data for LLM
@@
To facilitate research on table understanding and reasoning under challenging conditions, a number of datasets have been developed. Below, we introduce several representative datasets that are widely used for semi-structured table understanding.
@@
TEMPTABQA gupta2023temptabqa contains 11,454 question–answer pairs centered on temporal queries while SPREADSHEETBENCH ma2024spreadsheet provides a challenging benchmark for spreadsheet manipulation, consisting of 912 questions derived from real-world scenarios. MiMoTable li2024mimotable targets reasoning across multiple sheets and files, with 1,719 queries distributed over 428 spreadsheets. Evaluation results on these benchmarks reveal a significant performance gap, ranging from 20% to 50% between state-of-the-art models and human performance, underscoring the need for further advances in this area. INFOTABS gupta2020infotabs is designed for the natural language inference task on semi-structured tables. It contains 23,738 premise–hypothesis pairs, where premises are Wikipedia infoboxes and hypotheses are short sentences. WikiTableQuestions pasupat2015compositional focuses on question answering over semi-structured HTML tables, comprising 22,033 question–answer pairs together with their associated tables. FeTaQA nan2022fetaqa is a free-form table question answering dataset containing 10k Wikipedia-based pairs of tables, questions, answers, and supporting table cells. Unlike extractive QA datasets, FeTaQA provides human-written free-form answers including entities and their high-level relations. For comparison, Table II summarizes the characteristics of different datasets. The table highlights structural complexity including total cell counts and the number of merged cells, as well as question–answering types, including: (1) Content Match, where the task is to locate the correct cells, (2) Numeric Computation, which requires performing arithmetic or aggregation operations, and (3) Semantic Awareness, which involves semantically related or reasoning-intensive tasks. For instance, INFOTABS contains a large number of natural language inference questions, whereas SPREADSHEETBENCH features long, complex natural language queries (often exceeding 50 words) that describe multi-step tasks.
@@ 4.1 Chart
Takeaway: Chart is mainly a graph-based data presentation method to organize structured information into a form that is easy for human to observe. The future for LLM chart data analysis lies in designing a mechanism that mimics human attention for chart observation, whether it is end-to-end or pipeline-based reasoning.
@@
From a technical perspective, chart understanding poses unique challenges as it requires integrating visual perception with structured data extraction, numerical reasoning, and logical inference. As shown in Figure 7, tasks on charts show diverse features, requiring a multi-perspective understanding on the structure and content of charts. As a result, multiple tools are useful for chart understanding, each focusing on a specific side of the task. Research in this domain has primarily followed two paradigms: (1) pipeline-based methods, which decompose the task into modality conversion and semantic reasoning, and (2) end-to-end models, which directly map chart images and instructions to responses within a unified framework. In the following, we provide an overview of chart-related tasks and examine these two LLM-based technical paradigms.
Chart Captioning. Chart Captioning (or Summarization, Chart-to-Text) refers to a task on generating a descriptive caption for a given chart. The caption should contain key information or serve as a summary of the graph. Early studies primarily target common graphical features such as axes and legends. These features enable different charts can be summarized by successively mentioning all of the features, giving rise to planning-based pipelines for caption generation. For instance, Mittal et al mittal-etal-1998-describing propose feature-specific strategies for different types of graphs, while Reiter reiter-2007-architecture outline a four-stage data-to-text framework comprising (1) Signal Analysis to look for possible patterns and trends, (2) Data Interpretation to identify more complex messages, (3) Document Planning to decide which of the messages should be mentioned in the generated text, and (4) Microplanning and Realisation to create an actual text. Such pipeline-based approaches are stable and extensible in fixed scenarios but tend to be heavy and lack general applicability. Building on this line of work, ChartThinker liu2024chartthinker employs a large-scale dataset and a chain-of-thought strategy to enhance the logical coherence and accuracy of chart summarization. Leveraging LLMs, it first identifies the chart type and then iteratively generates reasoning steps to capture a holistic understanding of the chart.
@@ 4.2 Video
Takeaway: Through reformulating videos as structured token sequences that contains temporal information as well as visual evidence, LLM can track and understand specific events in object level, and perform downstream tasks
Video is inherently a sequence of frames whose spatial content evolves along the temporal dimension. Effective analysis therefore requires models to jointly capture spatial semantics and temporal dynamics wang2024grounded . Traditional approaches typically rely on computer vision backbones to extract frame-level representations, followed by aggregation mechanisms such as pooling, token merging, or attention to model temporal dependencies. While these methods provide reliable performance, the architectures are often manually engineered while the process is annotation-intensive and computationally inefficient for long video sequences, frequently sacrificing fine-grained temporal cues during compression.
Recent advances employ LLMs to enhance video understanding within multimodal frameworks. By representing videos as structured token sequences rather than isolated frames, LLM-based paradigms enable more coherent alignment between visual evidence and textual instructions. This shift addresses three major limitations of previous work: (1) computational inefficiency mitigated by adaptive frame selection and token reduction strategies; (2) loss of visual fidelity alleviated through selective filtering of redundant or irrelevant frames; and (3) weak temporal grounding improved by incorporating modules that explicitly encode temporal order and event duration shen2024tempme ; wang2024grounded . Rather than discarding existing vision encoders, lightweight adapters or temporal reasoning components are integrated to extend their capabilities while preserving efficiency.
@@
Temporally-Anchored. Existing video-language models often struggle with precise temporal localization, as it requires detecting sparse and transient signals within long and redundant video sequences. This involves (1) distinguishing relevant information from irrelevant frames (e.g., the first appearance or disappearance of a target object, the exact frames of a key gesture, the onset of a sound event) shen2024tempme , (2) identifying event boundaries (e.g., action start/end points, scene cuts, speech or audio segment boundaries), and (3) preserving semantic coherence over time (e.g., preventing association of an object’s attribute at $t_{1}$ with an unrelated action at $t_{5}$ while minimizing information loss and computational overhead) chen2024timemarker . Meanwhile, existing models face difficulties in handling videos of varying lengths, where fixed input sizes waste computation on redundant segments in short videos while uniform sampling misses critical events in longer ones.
@@
To address the above challenges, Temporally-Anchored methods explicitly align model predictions or reasoning with specific time points or intervals in a video. Concretely, these methods combine dynamic sampling (Adaptively select frames based on visual or semantic importance instead of uniformly sampling frames. For example, chen2024timemarker dynamically adjusts the sampling interval by monitoring motion intensity, allocating denser samples near action boundaries while sparsifying redundant segments.), temporal positional encodings (Models inject explicit time-aware representations into the token sequence. For instance, wang2024grounded augments frame embeddings with temporal position codes, enabling the model to reason about relative event order and duration.), and flexible tokenization (Including training-free similarity-based merging shen2024tempme , which groups redundant tokens based on feature similarity, and learnable saliency-driven merging lee2024video , which leverages gradient-based saliency to adaptively retain informative tokens.)
@@
Specifically, TimeMarker chen2024timemarker , a versatile Video-LLM based on LLaMA3-8B, is designed for high-quality video-grounded dialogue with temporal localization. The model introduces separator tokens when the scene changes to represent physical time and temporal separator tokens (like second 5.) to enhance temporal awareness. In addition, it incorporates an AnyLength mechanism for dynamic frame sampling and adaptive token merging, which adjusts the frames-per-second (FPS) and merges visual tokens to match the LLM’s context window and GPU budget. Similarly, Grounded-VideoLLM wang2024grounded augments the architecture with a temporal stream and discrete time tokens to enable explicit timestamp reasoning, trained through progressive video–language localization tasks. Seq2Time deng2025seq2time further extends this direction by adopting sequential knowledge transfer which learns temporal grounding in a self-supervised manner by converting ordered image or clip sequences into time-annotated training signals. Specifically, researchers use a 10-second clip from Kinetics-700 to serve as the fundamental unit. Multiple such clips are then concatenated into a sequence to simulate the progression of events in a longer video. Each clip is associated with a caption and its positional index within the sequence. The model is trained to predict or align the sequential order of these clips, thereby internalizing temporal order even without explicit timestamp labels.
@@
Instruction-Aware Relative Temporal Localization. In contrast to temporally-anchored approaches that output absolute timestamps or time-interval tokens (e.g., “at 00:02:15”), the instruction-aware relative temporal localization focuses on answering relative temporal queries (e.g., “what happened after the second bell?”).
Early efforts often integrate video foundation models (e.g., Conv3D/I3D-style models, two-stream models (models with architecture that process video stream through temporal stream and spatial stream separately) and transformer-based video encoders) with text-only LLMs, but such systems often required heavy pretraining and complex multi-module designs.
@@
RED-VILLM huang2024image demonstrates that existing image-based LLMs can be augmented with lightweight temporal adapters that align instruction semantics with temporal video features, enabling efficient instruction-conditioned reasoning without retraining a full video model. IVA-LLM li2024llms integrates a spatial feature interactor and a temporal frame selector within LLM blocks. When a query is asked by the user, it first utilizes the temporal frame selector to select the most relevant frames from the video based the user’s query. Once the temporal frame selector has identified the key frames,the spatial feature interactor allows the model to focus on specific visual regions or objects within a frame. Such architectures enhance efficiency and visual fidelity under token aggregation while maintaining strong temporal grounding aligned with natural language instructions.
@@
Video Emotional Analysis. Video emotional analysis seeks to infer affective states, social dynamics, and well-being indicators from multimodal cues. Understanding and promoting individual well-being is essential for staff health and long-term successmuller2024predicting . Current tools for assessing emotional changes rely on time-consuming surveys and questionnaires, which fails to find the subtle, context-dependent nature of emotion and the need for synchronized processing of visual, auditory, and textual signals.
@@
In order to provide the real-time feedback needed to raise awareness and change individual behavior, PERMA muller2024predicting aims to predict individual well-being during teamwork interactions. Specifically, they first record panoramic video using cameras. For each individual in the video, they compute nonverbal behavior features such as facial expression, head pose, 3D gaze direction, and environmental parameters like scene brightness. These are organized into time series and summarized into 125 statistical features. Finally, these features are aligned with self-reported PERMA survey responses, and machine learning models are trained to predict each PERMA pillar—treated as a classification or regression tasks to distinguish discrete well-being states Another system jadhav2023ai evaluates candidates during interviews by fusing face detection with HaarCascade classifiers(a face detection algorithm that identifies faces within an input image and provide the coordinates of each detected face adepu2020interviewee ), extracting audio features via Mel-frequency cepstral coefficients jadhav2023ai , and modeling visual and acoustic sequences using Inception-V3 and LSTM encoders. Textual data are pre-processed through standard NLP pipelines (tokenization, stop-word removal, possibly sentiment features) before integration, allowing the model to generate joint predictions of personality traits and performance from frame-level and clip-level inputs.
@@
Object Detection. Object detection localizes and classifies object instances in images by assigning bounding boxes and category labels. In video analysis, this task is further extended to ensure temporal consistency and to track specific objects across frames.

### v2

@@ 3.2 Semi-Structured Tables
This section focuses on semi-structured table processing, encompassing techniques for understanding, representing, and extracting information from such tables. Unlike structured relational data, semi-structured tables feature complex structures, including merged cells, hierarchical headers, and nested tables.
Semi-structured tables share a set of properties, which can be summarized into five core characteristics, illustrated in Figure 5: $(i)$ Wrong Index: row or column indices may be missing, ambiguous, or inconsistent suhara2022annotating . $(ii)$ Hierarchical Content: rows or columns may not follow a uniform schema (e.g., a “summary” cell under an “ID” column) yang2022tableformer . $(iii)$ Merged Cell: a single cell may span multiple rows or columns guo2022trust . $(iv)$ Flexible Header Orientation: keys and values can appear in varying positions (e.g., left-of or above their associated values). $(v)$ Inconsistent Content: cell values vary widely, including words, numbers, or full sentences.
@@
With the aid of LLMs, it becomes possible to directly reason over semi-structured tables, interpret their intricate layouts, and bridge the gap between natural language queries and heterogeneous table structures, thereby enabling more accurate and context-aware query answering.
Traditional table processing approaches like TAPAS herzig2020tapas mainly rely on lightweight models (e.g., BERTdevlin2019bert , RoBERTaliu2019roberta ) to analyze the table. These methods rely on specialized embedding mechanisms to capture the two-dimensional structure of tables, including rows, columns, and numerical order. To support tasks such as cell selection or aggregation, they usually require large-scale weakly supervised pre-training yang2022tableformer . However, due to the small number of parameters and weak NLP capabilities of lightweight models, this method cannot handle some tasks based on semi-structured tables with confusing layouts. Leveraging the strong semantic understanding capabilities of LLMs, LLM-based approaches generally preprocess the semi-structured table and then perform table understanding or reasoning. We conducted some research on this, and the work is shown in Figure 6.
Figure 6: Overview of LLM for Semi-Structured Table.
@@
$\bullet$ Tree-based Table Modeling. TUTA wang2021tuta introduces a tree-based encoding for semi-structured tables, where the root splits into row and column nodes, each expanding hierarchically into trees. Table entries are nodes with parents corresponding to their row and column (e.g., column nodes like “ID” or “Phone Number” and row nodes representing individual names). ST-Raptor tang2025st proposes a HO-Tree to encode tables, combining a Meta-Tree for header structure and semantics with a Body-Tree for content. It uses a VLM to detect meta-information, heuristic rules to segment tables, and a depth-first strategy to construct the HO-Tree.
$\bullet$ Model-based Structuring. TabFormer 10.1145/3726302.3730071 converts semi-structured tables into relational data using a chain-of-thought (CoT) approach. LLMs first apply step-by-step transformations with “soft operators” (e.g., splitting tables with multiple header groups). After preprocessing, SQL is used to reconstruct the table, including predicting indices, generating CREATE statements for the schema, and inserting the extracted contents into the relational table.
@@
Auto-Tables li2023autotablessynthesizingmultisteptransformations integrates OCR for structure recognition and applies domain-specific language (DSL) operators. These operators abstract common spreadsheet operations (e.g., melt, transpose), thereby enhancing generalization and adaptability in table conversion tasks. TradeSweep lee2024can , while closely related to Auto-Tables, differs in its approach. Instead of predefined DSL operators, it accepts user-specified transformation requirements and directly generates table conversion code by leveraging a code repository, offering greater flexibility in addressing diverse user needs.
Table Prompting. Table prompting converts table data into text input for LLMs. Traditional approaches, like table linearization, flatten all cell contents into a single string, often producing overly long prompts and losing structural information. To address this, table content and structure should be compressed, enabling LLMs to handle large tables while keeping structural loss controllable.
@@
SHEETCOMPRESSOR dong2024spreadsheetllm efficiently serializes large spreadsheets for LLM prompting. Structural Anchors identify key heterogeneous rows and columns, pruning redundant ones to create a skeletonized sheet. Inverted-Index Translation encodes non-empty cells in a JSON dictionary, merging duplicates to reduce storage. Data Format Aggregation clusters adjacent numerical cells of the same type or format, capturing distribution patterns without redundancy. HySem pp2024hysem introduces a lightweight LLM pipeline for local deployment. Its Context Optimizer compresses HTML table tokens so each cell uses only one or two words (e.g., compress “Theme 3: Factors affecting stuff response” into “Theme 3”), establishing a direct mapping between markup and content while minimizing token usage.
Table Querying. LLMs require effective methods to extract table contents, known as table querying. While conceptually similar to SQL, this section focuses on approaches that directly extract information from semi-structured tables rather than traditional SQL querying. CoS dong2024spreadsheetllm is a downstream table query system built on SHEETCOMPRESSOR. Given a spreadsheet and a natural language query, it first performs Table Identification and Boundary Detection to locate the relevant subtable. The subtable and query are then passed to the LLM for focused, context-aware response generation. In ST-Raptor tang2025st , table querying relies on the HO-Tree. Complex questions are decomposed into sub-questions mapped to sequences of basic operations (e.g., retrieval, condition). The alignment operation grounds abstract query terms to actual headers or values, while Top-down and Bottom-up retrieval navigate the HO-Tree. For large tables, Column-Type Annotation (e.g., discrete or continuous) limits the search space. A dual validation mechanism—forward constraints for intermediate consistency and backward reasoning to reconstruct the query—reduces errors and hallucinations.
@@
Semi-Structured Data for LLM To facilitate research on table understanding and reasoning under challenging conditions, a number of datasets have been developed. Below, we introduce several representative datasets that are widely used for semi-structured table understanding.
@@
TEMPTABQA gupta2023temptabqa contains 11,454 question–answer pairs centered on temporal queries while SPREADSHEETBENCH ma2024spreadsheet provides a challenging benchmark for spreadsheet manipulation, consisting of 912 questions derived from real-world scenarios. MiMoTable li2024mimotable targets reasoning across multiple sheets and files, with 1,719 queries distributed over 428 spreadsheets. Evaluation results on these benchmarks reveal a significant performance gap, ranging from 20% to 50% between state-of-the-art models and human performance, underscoring the need for further advances in this area. INFOTABS gupta2020infotabs is designed for the natural language inference task on semi-structured tables. It contains 23,738 premise–hypothesis pairs, where premises are Wikipedia infoboxes and hypotheses are short sentences. WikiTableQuestions pasupat2015compositional focuses on question answering over semi-structured HTML tables, comprising 22,033 question–answer pairs together with their associated tables. FeTaQA nan2022fetaqa is a free-form table question answering dataset containing 10k Wikipedia-based pairs of tables, questions, answers, and supporting table cells. Unlike extractive QA datasets, FeTaQA provides human-written free-form answers including entities and their high-level relations. For comparison, Table II summarizes the characteristics of different datasets. The table highlights structural complexity including total cell counts and the number of merged cells, as well as question–answering types, including: $(i)$ Content Match, where the task is to locate the correct cells, $(ii)$ Numeric Computation, which requires performing arithmetic or aggregation operations, and $(iii)$ Semantic Awareness, which involves semantically related or reasoning-intensive tasks. For instance, INFOTABS contains a large number of natural language inference questions, whereas SPREADSHEETBENCH features long, complex natural language queries (often exceeding 50 words) that describe multi-step tasks.
@@
Takeaway: Semi-structured data remains challenging due to its irregular and nested nature. LLMs show strong potential to address these difficulties through three main directions. Complex Data Understanding (O1): Enable reasoning over irregular and nested structures. NL-based Interfaces (O2): Facilitate intuitive interaction and query generation through natural language. Semantic Operators (O3): Empower models to perform semantic-level operations such as structural interpretation, summarization, and context-aware reasoning.
@@ 4.1 Chart
From a technical perspective, chart understanding poses unique challenges as it requires integrating visual perception with structured data extraction, numerical reasoning, and logical inference. As shown in Figure 7, tasks on charts show diverse features, requiring a multi-perspective understanding on the structure and content of charts. As a result, multiple tools are useful for chart understanding, each focusing on a specific side of the task. Research in this domain has primarily followed two paradigms: $(i)$ pipeline-based methods, which decompose the task into modality conversion and semantic reasoning, and $(ii)$ end-to-end models, which directly map chart images and instructions to responses within a unified framework. In the following, we provide an overview of chart-related tasks and examine these two LLM-based technical paradigms.
Chart Captioning. Chart Captioning (or Summarization, Chart-to-Text) refers to a task on generating a descriptive caption for a given chart. The caption should contain key information or serve as a summary of the graph. Early studies primarily target common graphical features such as axes and legends. These features enable different charts can be summarized by successively mentioning all of the features, giving rise to planning-based pipelines for caption generation. For instance, Mittal et al mittal-etal-1998-describing propose feature-specific strategies for different types of graphs, while Reiter reiter-2007-architecture outline a four-stage data-to-text framework comprising $(i)$ Signal Analysis to look for possible patterns and trends, $(ii)$ Data Interpretation to identify more complex messages, $(iii)$ Document Planning to decide which of the messages should be mentioned in the generated text, and $(iv)$ Microplanning and Realisation to create an actual text. Such pipeline-based approaches are stable and extensible in fixed scenarios but tend to be heavy and lack general applicability. Building on this line of work, ChartThinker liu2024chartthinker employs a large-scale dataset and a chain-of-thought strategy to enhance the logical coherence and accuracy of chart summarization. Leveraging LLMs, it first identifies the chart type and then iteratively generates reasoning steps to capture a holistic understanding of the chart.
@@
Takeaway: Charts present structured data in a visually interpretable form. Future work should focus on mechanisms that emulate human attention for chart understanding, whether through end-to-end or pipeline-based reasoning. Complex Data Understanding (O1): Process charts with different patterns and organize data results mittal-etal-1998-describing ; reiter-2007-architecture ; liu2024chartthinker ; masry2024chartgemmavisualinstructiontuningchart . NL-Based Interface (O2): Pay attention to details that users care about while allowing NL interactions das2025chartsofthoughtenhancingllmvisualization ; ye2024mplugowlmodularizationempowerslarge . Semantic Operators (O3): Capture semantic information hidden in positional relationships among the elements masry2023unichart ; wu2024chartinsights ; xu2025chartmoemixturediverselyaligned . Autonomous Evolution (O4): Notice the distribution of information in charts, and extend its ability to solve a lot of downstream taskshan2023chartllama ; huang2025evochart .
@@ 4.2 Video
Video inherently represents evolving spatial content over time, requiring models to jointly capture spatial semantics and temporal dynamics wang2024grounded . Traditional methods use vision backbones for frame-level features followed by pooling, token merging, or attention to model temporal relations. Although effective, these architectures are manually designed, annotation-heavy, and computationally inefficient for long sequences, often losing fine-grained temporal details during compression.
Recent advances integrate LLMs into multi-modal video understanding by representing videos as structured token sequences instead of independent frames. This design improves visual–text alignment and addresses key limitations of prior methods: $(i)$ computational inefficiency via adaptive frame selection and token reduction, $(ii)$ visual fidelity loss through selective filtering of redundant frames, and $(iii)$ weak temporal grounding using modules that encode event order and duration shen2024tempme ; wang2024grounded . Rather than replacing vision encoders, lightweight adapters and temporal reasoning components extend their capabilities while maintaining efficiency.
@@
Temporally-Anchored. Existing video–language models struggle with precise temporal localization, as it requires detecting sparse, transient signals within long and redundant sequences. This entails $(i)$ distinguish relevant information (e.g., object appearance, key gestures, or sound onsets) shen2024tempme , $(ii)$ identifying event boundaries (e.g., action start/end or scene transitions), and $(iii)$ maintaining temporal coherence by preventing misalignment of object states or actions across time chen2024timemarker . Moreover, handling videos of varying lengths remains challenging, where fixed input sizes waste computation on short clips, and uniform sampling in long videos often overlooks critical events.
@@
To tackle these challenges, Temporally-Anchored methods explicitly align model reasoning with specific time points or intervals. They integrate three key techniques: $(i)$ dynamic sampling, which adaptively selects frames by visual or semantic importance (e.g., chen2024timemarker adjusts sampling density based on motion intensity, allocating more frames near action boundaries); $(ii)$ temporal positional encoding, which injects explicit time-aware representations into token sequences (e.g., wang2024grounded augments frame embeddings with temporal position codes to model event order and duration); and $(iii)$ flexible tokenization, including training-free similarity-based merging shen2024tempme and learnable saliency-driven merging lee2024video , both designed to retain informative tokens while removing redundancy.
@@
Specifically, TimeMarker chen2024timemarker , a versatile Video-LLM based on LLaMA3-8B, enables high-quality video-grounded dialogue with temporal localization. It introduces separator tokens to represent physical time, temporal separator tokens to enhance temporal awareness, and an AnyLength mechanism for dynamic frame sampling and adaptive token merging which adjusts the frames-per-second (FPS) and visual tokens to match the LLM’s context window and GPU budget. Similarly, Grounded-VideoLLM wang2024grounded augments the architecture with a temporal stream and discrete time tokens to enable explicit timestamp reasoning, trained through progressive video–language localization tasks. Seq2Time deng2025seq2time advances this direction through sequential knowledge transfer, learning temporal grounding in a self-supervised manner by converting ordered image or clip sequences into time-annotated signals. Specifically, 10-second clips from Kinetics-700 are concatenated to simulate event progression in longer videos, each paired with a caption and positional index. The model learns to predict or align their sequential order, internalizing temporal structure without explicit timestamps.
@@
Instruction-Aware Relative Temporal Localization. In contrast to temporally-anchored approaches that output absolute timestamps or time-interval tokens (e.g., “at 00:02:15”), the instruction-aware relative temporal localization focuses on answering relative temporal queries (e.g., “what happened after the second bell?”). Early studies typically combined video foundation models (e.g., Conv3D/I3D-style, two-stream, or transformer-based encoders) with text-only LLMs, but these systems relied on heavy pretraining and complex multi-module architectures.
@@
RED-VILLM huang2024image shows that image-based LLMs can be efficiently extended to video understanding by introducing lightweight temporal adapters that align instruction semantics with temporal video features, avoiding full video-model retraining. Similarly, IVA-LLM li2024llms integrates a spatial feature interactor and a temporal frame selector within LLM blocks. The selector retrieves query-relevant frames, while the interactor focuses on salient regions or objects within those frames. This design enhances efficiency and visual fidelity under token aggregation while preserving strong temporal grounding aligned with natural language instructions.
@@
Video Emotional Analysis. Video emotion analysis aims to infer affective states, social dynamics, and well-being from multimodal cues. As understanding and promoting well-being is crucial for long-term health and performance muller2024predicting , traditional survey-based assessments are limited by their time-consuming nature and inability to capture the subtle, context-dependent, and synchronized dynamics of visual, auditory, and textual signals.
@@
To provide real-time feedback for behavior awareness, PERMA muller2024predicting predicts individual well-being during teamwork interactions. Panoramic videos are recorded, and for each person, nonverbal cues, including facial expressions, head pose, and 3D gaze, and environmental factors like scene brightness are extracted and summarized into 125 statistical features. These features are aligned with self-reported PERMA survey responses, and machine learning models predict each PERMA pillar as either classification or regression tasks, capturing discrete well-being states. Another system jadhav2023ai assesses interview candidates by fusing multimodal cues. Faces are detected using HaarCascade classifiers adepu2020interviewee , audio is represented via Mel-frequency cepstral coefficients, and visual and acoustic sequences are modeled with Inception-V3 and LSTM encoders. Textual data undergo standard NLP preprocessing before integration (e.g., tokenization, stop-word removal, sentiment features), enabling joint predictions of personality traits and performance from frame- and clip-level inputs.
@@
Object Detection. Object detection identifies and classifies instances in images by assigning bounding boxes and category labels. In video analysis, it extends to maintain temporal consistency and track objects across frames.
Recent methods isolate region-level tokens and maintain object-specific trajectories over time. To capture fine-grained spatial and temporal details, VideoRefer yuan2025videorefer uses a spatial–temporal object encoder with adaptive temporal token merging to generate precise region representations for single- and multi-frame reasoning.